OpenStudy (anonymous):

Solve e^(4-x)=(e^2)*(e^(x^(-2)))

sam (.sam.):

is it\[\Large e^{4-x}=e^2 e^{\frac{1}{x^2}}\]

OpenStudy (anonymous):

yes